# Topology and Algebra of Bonded Knots and Braids

### v1

@@ Abstract.
bonded links, bonded braids, bonded bracket, enhanced bonds, bonded knotoids, bonded braidoids, L-move, Alexander theorem, Markov theorem.
@@ 0. Introduction
In this paper, we develop the theory of bonded knots, braids, knotoids and braidoids. For illustrations of these objects see Figures 1, 21, 39, 74, and 81. These structures reflect different physical situations which can occur in applications, such as protein folding and molecular biology. We consider three types of bonds: long bonds which can be knotted or linked and are not local, standard bonds which are in the form of straight segments and do not cross between themselves, and tight bonds which occur locally and are nearly equivalent to graphical nodes in the mathematical formalism. We also discuss two kinds of isotopy for the three categories of bonded knots and links, long, standard, and tight, which reflect different physical assumptions about the flexibility of bonds: topological vertex isotopy, whereby the nodes of the bonds can move freely, and rigid vertex isotopy, whereby the nodes of the bonds move along with rigid 3-balls in which they are embedded.
It is a theme of this paper to compare long bonds, standard bonds and tight bonds in the topological and rigid categories. We present the set of allowed moves for each category and highlight forbidden moves that distinguish bonded links from related concepts like tied links. We also recall and extend invariants for bonded knots: notably, the unplugging technique for the topological vertex category, the tangle insertion for the rigid vertex category, and a Kauffman bracket type polynomial (the bonded bracket) that is invariant under regular isotopy of rigid vertex bonded links.
@@
While knotoids address the issue of open endpoints, real biomolecules such as proteins often have additional internal connections that influence their topology. For instance, many proteins form bonds such as disulfide bridges, hydrogen bonds, or salt bridges, which link distant parts of the chain. These bonds can create topologically nontrivial structures (e.g. forming loops or lassos) even if the protein backbone is not closed. A classical knot or knotoid representation of a protein does not capture the presence of such bonds. To incorporate these features, Kauffman, Magarshak [20] and others [30] developed new knot invariants that take into account “stem” regions or paired interactions in biomolecules. In particular, studies of RNA tertiary structure and protein folding in the 1990s and 2000s led to polynomial invariants sensitive to base pair bonds or disulfide bonds.
@@
These developments eventually led to the notion of a bonded knotoid [11] and more recently of a bonded link, which is a knotoid or link augmented by embedded arcs representing bonds between points on the strands. Bonded links were formalized as topological models for protein chains with bridges in [30, 11] and subsequent works [9]. In a bonded link, the usual knot or link diagram is supplemented with one or more distinguished arcs, called bonds, connecting points on the knot/link strands (the bonds are drawn as, say, orange arcs in diagrams). These bonds represent physical connections that are not part of the covalently linked backbone. For example, a disulfide bond in a protein can be modeled as a bond connecting two cysteine points on the chain.
@@
Mathematically, a bond is an embedded arc whose endpoints lie on (distinct or possibly the same) components of the link diagram. The incorporation of bonds into knot diagrams gives rise to a rich extension of knot theory, with new moves and new invariants. A number of recent papers have studied invariants of bonded links and knotoids, see for example [20, 11, 9]; In this paper an extended bracket polynomial invariant is defined for bonded links. Invariants of bonded knotoids have been applied to protein structure classification, providing a novel way to distinguish different folding patterns of open proteins [11, 9].
@@
We note that the interaction lines in Feynman diagrams have the formal structure of bonds in our sense. Thus a Feynman diagram can be regarded as a bonded graph (possibly with a knotted embedding in three space). In fact, just such ideas are in back of the work in the book “Knots and Feynman Diagrams” by Dirk Kreimer [23] where the bonds in the Feynman diagrams undergo tangle insertion (in our sense) and are thereby associated with specific knots and links. Kreimer suggests that the topological types of these knots and links associated with the diagrams are significantly related to the physical evaluations of the diagrams.
@@
In Section 1, we develop the notion of long bonded links and we establish the allowed isotopy moves in both topological and rigid vertex categories. In Section 2, we introduce standard bonded links, namely, long bonded links with unknotted and unlinked bonds, and we establish that any long bonded link can be isotoped into standard form. We also give a full set of isotopy moves for standard bonded links in both topological and rigid vertex categories. In Section 3, we introduce a stricter category of bonded links, called tight bonded links. Tight bonded links are standard bonded links whose bonds do not interact with link arcs. We also present a set of isotopy moves for this setting.
@@
In Section 5, we turn to the algebraic counterpart of bonded links: bonded braids. We define bonded braids as braids with bonds connecting strands, and we introduce the bonded braid monoid $BB_{n}$ on $n$ strands, giving a complete presentation by generators and relations and two reduced ones. Moreover, we establish its isomorphism to the singular braid monoid. In Section 6 we prove the analogue of the Alexander theorem in this setting: every bonded link can be obtained as the closure of a bonded braid (Alexander theorem for bonded links).
In Section 7 we prove bonded braid equivalence theorems, showing that two bonded braids have equivalent closures if and only if they are related by certain moves (Markov moves adapted to bonds). We present an $L$-move formulation of braid equivalence, articulating the Markov theorem by using local cutting and rejoining moves ($L$-moves) on bonded braids. This provides a more fundamental understanding of how bond isotopies translate into algebraic moves. Finally, we identify some forbidden moves due to bonds.
In Section 8, we introduce enhanced bonded links and braids. An enhanced bond comes in two types, representing, for example, an attractive vs. a repulsive interaction. Mathematically, an attracting bond and a repelling bond may be thought of as mutual inverses. We show how allowing two bond types effectively turns the bonded braid monoid into a bonded braid group, and we define this enhanced bonded braid group $EB_{n}$ with its extended generator set. The analogues of the Alexander and the Markov theorems are established for the enhanced setting as well.
@@
We conclude with a discussion of further directions, including the formulation of a bonded Morse category and the extension of the braidoid–knotoid correspondence to the bonded and enhanced bonded settings, which we plan to develop in future work. This work was developed independently of work by others, such as [9].
@@ 1.1. Definition of Bonded Links
A bonded link (or bonded knot, in case of a single component) is a pair $(L,B)$, where $L$ is an (oriented) link in $S^{3}$, and $B$ is a set of $k$ disjoint arcs properly embedded in the complement $S^{3}\setminus L$, such that each bond arc in $B$ has its two endpoints, called nodes, attached on $L$. The nodes are attached transversely to $L$, with each node attaching at a distinct point on $L$. If $B=\emptyset$, then $(L,\emptyset)$ is just a classical link.
@@ 1.2. Bonded Isotopy and Allowed Moves
Clearly, rigid vertex isotopy implies topological vertex isotopy. On the level of diagrams, this means that one can go from a diagram of $(L_{1},B_{1})$ to a diagram of $(L_{2},B_{2})$ by a finite sequence of planar isotopies and a set of local Reidemeister moves involving any types of arcs (link arcs or bonds or both) respecting or not rigid vertex isotopies. To see this, we adapt [18, Theorem 2.1] and [18, Section III] to the bonded link setting and we obtain the following:
@@
Similarly, two (oriented) bonded links $(L_{1},B_{1})$ and $(L_{2},B_{2})$ are equivalent as bonded links via rigid vertex isotopy if corresponding diagrams of theirs differ by a finite sequence of the moves (i)-(v) above together with the rigid vertex twists (RVT) shown in Figure 8, excluding the TVT moves (Figure 7). Rigid vertex isotopy moves are illustrated in Figures 2, 3, 4, 5, 6, and 8.
@@
Figure 2. The planar isotopy moves
@@
Proof of Proposition 1.4. This is an adaptation of [18, Theorem 2.1] to the case of trivalent topological graphs in the particular form of bonded links. The only differences are as follows. Kauffman’s original statement of the basic moves of topological vertex isotopy on embedded graphs includes the following moves (adapted to the bonded category) in place of TVT and VS:
@@
These moves, however, follow easily from the moves in Figures 3, 6, and 7, as explicitly demonstrated in Figures 10 and 11.
@@
The case of rigid vertex isotopy for bonded links is an adaptation of [18, Section III] to the case of trivalent rigid vertex graphs in the particular form of bonded links and follows analogously. Hence the proof is complete. $\Box$
@@ 1.3. More Diagrammatic Moves for Bonded Links
We now present additional local diagrammatic moves in the setting of topological bonded links, and show that they all follow from the list of basic moves stated in Proposition 1.4 and exemplified in Figures 2345, and 6.
@@
Bond–crossing interaction. Next, we consider moves involving the interaction of a bond with a crossing. Three scenarios arise: (i) both arcs of the crossing lie within the region of a bond, (ii) one arc provides a bonding site, (iii) both arcs provide a bonding site. The first two comprise the crossing slide moves (Figures 15 and 16); the latter gives rise to the bonded flype and bonded double flype moves (Figure 18). In Figures 15, 16, 17 and 18 we show that these moves follow from the basic topological moves listed in Proposition 1.4.
@@ 2. Standard Bonded Links
Two (oriented) standard bonded links are topologically isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 3 and 23 with all variants.
Similarly, two (oriented) standard bonded links are rigid vertex isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures  3 and 24 with all variants.
@@ 3. Tight Bonded Links
Two (oriented) tight bonded links are topologically isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 3 and 28. Similarly, two (oriented) tight bonded links are rigid vertex isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 3 and 29.
@@ 4. Invariants of Topological and Rigid Vertex Bonded Links
4.1. Topological Bonded Links: The unplugging operation
For a bonded link $(K,B)$, we view the link together with the bonds as a trivalent embedded graph. At each node (vertex) where a bond attaches, we may perform an unplugging operation: removing one of the edges incident to the vertex. This yields a collection of classical link diagrams corresponding to the various choices of unpluggings. See Figure 31.
@@ 4.1. Topological Bonded Links: The unplugging operation
A classical result due to Kauffman asserts:
Theorem 4.1 (Kauffman).
Let $(K,B)$ be a bonded link. Then, considering the bonds and the link arcs as graph edges invariantly:
  • Consequently, any ambient isotopy invariant applied to $\mathcal{L}_{(K,B)}$ is also a topological vertex isotopy invariant of $(K,B)$.

### v2

@@ Abstract.
bonded links, long bonds, standard bonds, tight bonds, topological vertex equivalence, rigid vertex equivalence, unplugging, tangle insertion, bonded bracket polynomial, bonded braids, bonded braid monoid, enhanced bonds, bonded braid group, bonded knotoids, bonded braidoids, $L$-moves, Alexander theorem, Markov theorem.
@@ 0. Introduction
In this paper, we develop the theory of bonded knots, braids, knotoids and braidoids. For illustrations of these objects see Figures 1 and 2. These structures reflect different physical situations which can occur in applications, such as protein folding and molecular biology. We consider three types of bonds: long bonds which can be knotted or linked and are not local, standard bonds which are in the form of straight segments and do not cross between themselves, and tight bonds which occur locally and are nearly equivalent to graphical nodes in the mathematical formalism. We also discuss two kinds of isotopy for the three categories of bonded knots and links, long, standard, and tight, which reflect different physical assumptions about the flexibility of bonds: topological vertex isotopy, whereby the nodes of the bonds can move freely, and rigid vertex isotopy, whereby the nodes of the bonds move along with rigid 3-balls in which they are embedded.
It is a theme of this paper to compare long bonds, standard bonds and tight bonds in the topological and rigid categories. We present the set of allowed moves for each category and highlight forbidden moves that distinguish bonded links from related concepts like tied links. We also recall and extend invariants for bonded knots: notably, the unplugging technique for the topological vertex category, the tangle insertion for the rigid vertex category, and a Kauffman bracket type polynomial (the bonded bracket) that is invariant under regular isotopy of rigid vertex bonded links. See [24, 28]. Continuing this theme we introduce and study bonded braids as the algebraic counterpart of bonded links. We establish the bonded braid monoid, we point out its relation to the singular braid monoid [4] and we extend it to the bonded braid group. We further establish the interaction between topological bonded links and bonded braids by means of a closure operation, a braiding algorithm and bonded braidoid equivalences. This algebraic structure could be used in encoding the topological structures of bonded links and of objects that they model.
Figure 1. On the left: a bonded link. On the right: a bonded braid and its closure.
@@
Figure 2. On the left: a bonded knotoid. On the right: a bonded braidoid and its closure
@@
While knotoids address the issue of open endpoints, real biomolecules such as proteins often have additional internal connections that influence their topology. For instance, proteins form bonds such as disulfide bridges, hydrogen bonds, or salt bridges, which link distant parts of the chain. These bonds can create topologically non-trivial structures (e.g. forming loops or lassos) even if the protein backbone is not closed [29, 39, 36, 11]. A classical knot or knotoid representation of a protein does not capture the presence of such bonds. To incorporate these features, Kauffman, Magarshak [26] and others [40] developed new knot invariants that take into account “stem” regions or paired interactions in biomolecules. In particular, studies of RNA tertiary structure and protein folding in the 1990s and 2000s led to polynomial invariants sensitive to base pair bonds or disulfide bonds.
@@
These developments eventually led to the notion of a bonded knotoid [15] and more recently of a bonded link, which is a knotoid or link augmented by embedded arcs representing bonds between points on the strands. Bonded knotoids or links were formalized as topological models for protein chains with bridges in [40, 15] and subsequent works [1, 13]. In a bonded link or knotoid, the usual link or knotoid diagram is supplemented with one or more distinguished arcs, called bonds, connecting points on the knotoid/link strands (the bonds are drawn as, say, orange arcs in diagrams). These bonds represent physical connections that are not part of the covalently linked backbone. For example, a disulfide bond in a protein can be modelled as a bond connecting two cysteine points on the chain [11].
@@
Mathematically, a bond is an embedded arc whose endpoints lie on (distinct or possibly the same) components of the link diagram. The incorporation of bonds into knot diagrams gives rise to a rich extension of knot theory, with new moves and new invariants. A number of recent papers have studied invariants of bonded links and knotoids, with applications to protein structure classification, providing a novel way to distinguish different folding patterns of open proteins, see for example [26, 15, 1, 13].
In this paper we consider bonded multi-knotoids in full generality, where the bonds can be knotted and linked, and also up to both equivalence relations: rigid vertex as well as topological vertex equivalence. We also introduce and study bonded braidoids and their interrelation with bonded multi-knotoids.
@@
In regard to interdisciplinary connections, we note that the interaction lines in Feynman diagrams have the formal structure of bonds in our sense. Thus a Feynman diagram can be regarded as a bonded graph (possibly with a knotted embedding in three space). In fact, just such ideas are in back of the work of Kreimer in the book “Knots and Feynman Diagrams” [31] where the bonds in the Feynman diagrams undergo tangle insertion (in our sense) and are thereby associated with specific knots and links. Kreimer suggests that the topological types of these knots and links associated with the diagrams are significantly related to the physical evaluations of the diagrams.
In another direction, we note that current suggestions about knotted glueballs (closed loops of gluon flux related to the structure of protons) can be seen in our context as knotted structures consisting entirely of bonds. The strings of gluon field form highly attracting bonds between quarks in this model [38, 7]. Finally, we point out that there is ongoing research in the interface of molecular biology and the construction of molecules with specified polyhedral shapes [23]. This research also involves bonds to which our modelling applies.
@@
In Section 1, we develop the notion of long bonded links and we establish the allowed isotopy moves in both topological and rigid vertex categories. In Section 2, we introduce standard bonded links, namely, long bonded links with unknotted and unlinked bonds, and we establish that any long bonded link can be isotoped into standard form. We also give a full set of isotopy moves for standard bonded links in both topological and rigid vertex categories. We also identify some forbidden moves due to bonds. In Section 3, we introduce a stricter category of bonded links, called tight bonded links. Tight bonded links are standard bonded links whose bonds do not interact with link arcs. We also present a set of isotopy moves for this setting.
@@
In Section 5, we turn to the algebraic counterpart of bonded links: bonded braids. We define bonded braids as braids with bonds connecting strands, and we introduce the bonded braid monoid $BB_{n}$ on $n$ strands, giving a complete presentation by generators and relations and two reduced ones. Moreover, we establish its isomorphism to the singular braid monoid. In Section  6 we prove the analogue of the Alexander theorem in the topological bonded setting: every topological bonded link can be obtained as the closure of a bonded braid.
In Section 7 we prove bonded braid equivalence theorems, showing that two bonded braids have topological equivalent closures if and only if they are related by certain moves (Markov moves adapted to bonds) and bonded braid isotopy. We begin with an $L$-move formulation of bonded braid equivalence, by using local cutting and rejoining moves ($L$-moves) on bonded braids. This provides a more fundamental understanding of how bond isotopies translate into more algebraic moves analogous to the classical Markov theorem.
In Section 8, we introduce enhanced bonded links and braids. An enhanced bond comes in two types, representing, for example, an attractive vs. a repelling interaction. Mathematically, an attracting bond and a repelling bond may be thought of as mutual inverses. We show how allowing two bond types effectively turns the bonded braid monoid into a bonded braid group, and we define this enhanced bonded braid group $EB_{n}$ with its extended generator set. The analogues of the Alexander and the Markov theorems are established for the enhanced setting as well.
@@
We conclude with a discussion of further directions, including the formulation of a bonded Morse category and the extension of the braidoid–knotoid interaction to the bonded and enhanced bonded settings, which we plan to develop in future work.
Much of this work has been presented in conferences in a number of places. Here is a list of some of these presentations:
SL: (Topological) bonded knots and bonded knotoids; Knots and Proteins Workshop in Ljubljana, Slovenia, 12–14 February 2024 & Quantum Topology Seminar, SUSTech, 10 April 2024.
LK: Topological Bonded Knots and Knotoids; Banff International Research Station, March 21, 2024.
SL: Braids, braidings and braid equivalences; SKCM2 WPI Hiroshima U Winter School, 9–20 December 2024.
SL: Bonded knots, bonded knotoids and bonded DP tangles; SKCM2 Hiroshima U Winter School, 9–20 December 2024.
ID: Bonded Knots and Braids: From Topology to Algebra and Protein Structure; Applied Topology Day, Vrije Universiteit Amsterdam, April 8, 2025.
ID: Bonded knots and bonded braids; Algebraic and Geometric methods of Analysis, AGMA2025, May 26–29, 2025, Ukraine.
@@ 1.1. Definition of Bonded Links
A bonded link (or bonded knot, in case of a single component) is a pair $(L,B)$, where $L$ is a link in $S^{3}$, and $B$ is a set of $k$ disjoint arcs properly embedded in the complement $S^{3}\setminus L$, such that each bond arc in $B$ has its two endpoints, called nodes, attached on $L$. The nodes are attached transversely to $L$, with each node attaching at a distinct point on $L$. If $B=\emptyset$, then $(L,\emptyset)$ is just a classical link.
@@ 1.2. Bonded Isotopy and Allowed Moves
We shall further define the notion of trivial bonded link, which in knot theory corresponds to the notion of unknot and unlinks.
Definition 1.4.
A bonded link is trivial as topological resp. rigid bonded link if it can be turned by topological vertex isotopy resp. rigid vertex isotopy into a planar graph.
Figure illustrates a bonded knot which trivial as topological but non-trivial as rigid bonded knot. It also illustrates a trivial bonded link and a bonded link which is non trivial in either setting.
Clearly, rigid vertex isotopy implies topological vertex isotopy. On the level of diagrams, this means that one can go from a diagram of $(L_{1},B_{1})$ to a diagram of $(L_{2},B_{2})$ by a finite sequence of planar isotopies and a set of local Reidemeister moves involving any types of arcs (link arcs or bonds or both) respecting or not rigid vertex isotopies. To see this, we adapt [24, Theorem 2.1] and [24, Section III] to the bonded link setting and we obtain the following:
@@
Similarly, two (oriented) bonded links $(L_{1},B_{1})$ and $(L_{2},B_{2})$ are equivalent as bonded links via rigid vertex isotopy if corresponding diagrams of theirs differ by a finite sequence of the moves (i)-(v) above together with the rigid vertex twists (RVT) shown in Figure 10, excluding the TVT moves (Figure 9). Rigid vertex isotopy moves are illustrated in Figures 4, 5, 6, 7, 8, and 10.
@@
Figure 4. The planar isotopy moves.
@@
Proof of Proposition 1.5. This is an adaptation of [24, Theorem 2.1] to the case of trivalent topological graphs in the particular form of bonded links. The only differences are as follows. Kauffman’s original statement of the basic moves of topological vertex isotopy on embedded graphs includes the following moves (adapted to the bonded category) in place of TVT and VS:
@@
These moves, however, follow easily from the moves in Figures 5, 8, and 9, as explicitly demonstrated in Figures 12 and 13.
@@
The case of rigid vertex isotopy for bonded links is an adaptation of [24, Section III] to the case of trivalent rigid vertex graphs in the particular form of bonded links and follows analogously. Hence the proof is complete. $\Box$
@@ 1.3. More Diagrammatic Moves for Bonded Links
We now present additional local diagrammatic moves in the setting of topological bonded links, and show that they all follow from the list of basic moves stated in Proposition 1.5 and exemplified in Figures 4567, and 8.
@@
Bond–crossing interaction. Next, we consider moves involving the interaction of a bond with a crossing. Three scenarios arise: (i) both arcs of the crossing lie within the region of a bond, (ii) one arc provides a bonding site, (iii) both arcs provide a bonding site. The first two comprise the crossing slide moves (Figures 17 and 18); the latter gives rise to the bonded flype and bonded double flype moves (Figure 20). In Figures 17, 18, 19 and 20 we show that these moves follow from the basic topological moves listed in Proposition 1.5.
@@ 2. Standard Bonded Links
Two (oriented) standard bonded links are topologically isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 5 and 25 with all variants.
Similarly, two (oriented) standard bonded links are rigid vertex isotopic if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures  5 and 26 with all variants.
@@ 3. Tight Bonded Links
Two (oriented) tight bonded links are topologically isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 5 and 30. Similarly, two (oriented) tight bonded links are rigid vertex isotopic in the tight category if and only if any corresponding diagrams of theirs differ by a finite sequence of the moves illustrated in Figures 5 and 31.
@@ 4. Invariants of Topological and Rigid Vertex Bonded Links
4.1. Topological Bonded Links: Unplugging Operations
For a bonded link $(K,B)$, we view the link together with the bonds as a trivalent embedded graph. At each node (vertex) where a bond attaches, we may perform an unplugging operation: removing one of the edges incident to the vertex (view top row of Figure 33). At a 3-valent vertex there are three such choices. This yields a collection of unknotted (redundant) arcs and classical links corresponding to the various choices of unpluggings. A classical result due to Kauffman [24] asserts, in the language of bonded links:
Theorem 4.1.
Let $(K,B)$ be a bonded link. Then, considering the bonds and the link arcs as graph edges invariantly, we obtain:
[…]
Figure 33 bottom row shows the set of three knots derived by the unplugging of the given bonded knot, seen as an embedded graph.
@@ 4.1. Topological Bonded Links: Unplugging Operations
We note that we can create finer unplugging invariants for bonded links by restricting unplugging in relation to the bonds. There are two types, the bonded unplugging and the strict unplugging, which are complementary:
Definition 4.2.
The bonded unplugging of a bonded link $(K,B)$ is the unplugging where we simply unplug only the bonds, so we are left with the underlying link $K$.
It now follows immediately from Theorem 4.1 that:
Proposition 4.3.
[…]
The bonded unplugging can distinguish bonded knots. In Figure 34 the bonded unplugging yields an unknot in the one case and a trefoil knot in the other case. Hence the two bonded knots are distinguished. In this example with the knotted bond we have an embedding of a $\Theta$-graph. This knotted graph is the Turaev closure of the knotoid depicted by the knotted bond as a diagram in the surface of the 2-dimensional sphere. The fact that we have proved that the $\Theta$-graph has a non-trivial embedding implies that this knotoid is non trivial.
Figure 34. The bonded unplugging distinguishes the two bonded knots.
Definition 4.4.
The strict unplugging of a bonded link $(K,B)$ is the unplugging where at a given node only the arcs from the link are allowed to be unplugged.
At a 3-valent vertex there are two such choices. It also follows immediately from Theorem 4.1 that:
Proposition 4.5.
The resulting collection of knots and links (ignoring the unknotted arcs) from the strict unplugging of a bonded link is a topological vertex isotopy invariant of the bonded link.
Figure 35 illustrates an example of a bonded knot $G$ which is trivialized by the bonded unplugging. However, via the strict unplugging we obtain an unknot and a trefoil knot. This proves that the given bonded knot is non-trivial. Also that the $\Theta$-graph where one considers the bond as another edge is non-trivial. Let further $G_{u}$ be the same graph where now the upper arc in $G$ is a bond (see left illustration in Figure 35). If we perform bonded unplugging to $G_{u}$ we obtain a trefoil knot, therefore $G_{u}$ is not isotopic to $G$.
Figure 35. Unplugging edges from an embedded graph.
In the case of $G_{l}$ where the bond is the lower arc of $G$, the unplugging invariants do not distinguish $G_{l}$ and $G$. In fact $G_{l}$ and $G$ are isotopic as demonstrated in Figure 36.
Figure 36. Two isotopic bonded knots.